how do i solve 3/4=x/x+3 please show the work the book says that x=13 i cannot figure out how show the work
Answer by jim_thompson5910(35256)

3/4 = x/(x+3)

3(x+3) = 4x

3x+9=4x

3x=4x-9

3x-4x=-9

-x=-9

x=(-9)/(-1)

x=9

So the answer is x = 9. So there must be a typo somewhere.